A sports car accelerates from rest to 26.8 m/s (roughly 60 mi/h) in 5.1 seconds. What is the acceleration of
the car?
1
Expert's answer
2020-11-02T09:24:30-0500

By the definition of the acceleration, we get:


"a=\dfrac{v-v_0}{t}=\dfrac{26.8\ \dfrac{m}{s} -0\ \dfrac{m}{s}}{5.1\ s}=5.25\ \dfrac{m}{s^2}."

Answer:

"a=5.25\ \dfrac{m}{s^2}."